OpenStudy (hpfan101):

I'm having a hard time understanding partial fraction decomposition. \[\int\limits_{1}^{2}\frac{ x^3 + 4x^2 +x -1 }{ x^3 + x^2 }dx\]

OpenStudy (freckles):

if degs match or if degree is greater on top you must do division first

OpenStudy (hpfan101):

Ok so after I get the division I got: \[1+\frac{ 3x^2+x-1 }{ x^3+x^2 }\]

OpenStudy (hpfan101):

Did*

OpenStudy (freckles):

ok partial fraction time first step factor denominator

OpenStudy (hpfan101):

Ok, I know it can be factored into: (x^2)(x + 1) Not sure if I can factor it any more than that

OpenStudy (freckles):

ok cool and no

OpenStudy (freckles):

so x^2 is a linear factor^2 and x+1 is just a linear factor non-repeated

OpenStudy (freckles):

\[\frac{A}{x}+\frac{B}{x^2}+\frac{C}{x+1}=\frac{3x^2+x-1}{x^3+x^2}\]

OpenStudy (hpfan101):

Oh ok. I wasn't sure we were supposed to have A/x and then B/x^2. I thought there wouldn't be a coefficient over x.

OpenStudy (freckles):

I always do constant/linear^n 1 degree less than the function inside the power on the bottom

OpenStudy (hpfan101):

Ok that makes sense! So afterwards we would multiply everything by the denominator?

OpenStudy (freckles):

well we find the lcm which is x^2(x+1) use this to combine the fractions

OpenStudy (freckles):

example that first fraction I wrote we multiply top and bottom by x(x+1) second fraction we do (x+1) last fraction x^2

OpenStudy (freckles):

\[\frac{Ax(x+1)+B(x+1)+Cx^2}{x^2(x+1)}=\frac{3x^2+x-1}{x^3+x^2} \\ \implies Ax(x+1)+B(x+1)+Cx^2=3x^2+x-1 \\ \text{ now I have started \to really like heaviside method for this }\]

OpenStudy (hpfan101):

Ok I understand that part. Do we distribute next?

OpenStudy (freckles):

you can ... or use heaviside method...( I think you will like this more)

OpenStudy (freckles):

Since this is an equation that means we want both sides to be the same for any x

OpenStudy (freckles):

this means this equation should hold that is left side should equal right side when x=0 or when x=-1 ....and so on...

OpenStudy (freckles):

So enter in 0 for x on both sides you should get an easy equation to obtain B

OpenStudy (hpfan101):

Oh so B would end up being equal to -1?

OpenStudy (freckles):

yep now select x=-1 I'm choosing x=-1 because a lot of that stuff on the left hand side will disappear

OpenStudy (hpfan101):

Ok so c=1?

OpenStudy (freckles):

yep ok... now choose any value for x (an easy one preferably) and input your values for B and C into that equation and solve for A

OpenStudy (freckles):

any x you haven't used already*

OpenStudy (freckles):

like x=1

OpenStudy (freckles):

it doesn't matter which value for x you choose those you should wind up with A=2 (if I didn't make a mistake lol)

OpenStudy (hpfan101):

Yeah I got A=2 too when I chose x=2

OpenStudy (freckles):

I have to get ready to go out but I hope that helps @IrishBoy123 mentioning you because I have to leave

OpenStudy (hpfan101):

Ok thank you for the help! This way is much better!

OpenStudy (irishboy123):

good night @freckles !